Billy is 44 and she is twice as old as nina was when billy was as old as Nina is now. How old is nina now?
Answer by AnlytcPhil(1806)

Billy is 44, which is twice as old as Nina was x years ago, which was N-x

44 = 2(N-x)

x years ago, Billy, who is now 44, was 44-x, which is as old as Nina is now, N. 

44-x = N

The system of equations is

44 = 2(N-x)
44-x = N

Solve that system by substitution and get x = 11, N = 33

Nina is 33.

Checking: 11 years ago Billy was 44-11=33, which is Nina's present
age, and Nina was 33-11=22.  So Billy, at 44, is now twice as old as
Nina was 11 years ago at age 22.

Edwin